# Affective Computing in the Era of Large Language Models: A Survey from the NLP Perspective

### v1

@@ 4.4 Challenge
  • Despite their known universality, LLMs often struggle to simultaneously excel in both AU and AG within the domain of AC. Developing a unified model that effectively addresses AU and AG remains a significant challenge.
@@ 5 Prompt Engineering
LLMs, owning powerful understanding and generation capabilities, pave new ways for tasks associated with AC. Despite pre-training or supervised fine-tuning (SFT), effectively utilizing LLMs for AC remains challenging. Prompt engineering [27] achieves higher accuracy and reliability in the AC tasks by designing appropriate input prompts that guide LLMs to generate the desired outputs. This section introduces the latest advancements in prompt engineering for AC, focusing on key technologies and application cases to provide references and insights for related research. Current research identifies four main prompting techniques, as shown in Figure 3: Zero-shot Prompt, Few-shot Prompt, Chain-of-Thought (CoT) Prompt, and Agent.
Zero-shot Prompt [27] is a method where the model generates text about a specific topic or domain without any demonstrations or specific instances from training data. This approach is structurally simple, easy to understand, and capable of quickly adapting to different fields and tasks. By leveraging model’s extensive knowledge and intrinsic reasoning abilities, zero-shot prompts can effectively generate the required text. This makes them particularly effective in applications where high-quality content needs to be generated quickly, especially when data is insufficient or difficult to obtain.
Few-shot Prompt [27] provides models with a few examples or demonstrations related to the given task. These examples and the task description, are input as prompts to help the model better understand the specific task. The model can perform efficient reasoning and generate high-quality output by learning from a few examples and following the task instructions. Few-shot prompt improves the performance of many complex tasks and is particularly advantageous when there are limited data samples or data imbalances.
Figure 3: The Overview of Prompt Engineering
CoT Prompt [147] is a technique that significantly enhances the reasoning ability of LLMs. Instead of directly prompting the LLMs to generate output, the CoT prompting method uses a step-by-step reasoning process to guide the models through logical thinking. This method involves explicitly outputting intermediate steps as part of the reasoning chain. CoT prompting allows LLMs to generate more accurate and contextually relevant responses. It is effective in various complex reasoning tasks, such as arithmetic, common sense, and symbolic reasoning. Additionally, the intermediate steps make it easier for users to understand the thinking process of a model, thus improving the interpretability of the reasoning of LLMs.
Agent [92] is an emerging and promising prompt engineering technology. It functions as an intelligent agent capable of perceiving its environment, understanding it autonomously, making decisions, and executing actions. It can independently think and call upon various tools to gradually achieve goals. In an LLM Agent, the LLM acts as the agent’s brain, collaborating with multiple LLM components to complete each part of a task. The profile, tool, module, plan, and action modules perform their respective tasks based on the prompt method. The LLM brain then synthesizes the results from each module to achieve better outcomes. This collaborative approach makes agents particularly effective in solving complex tasks that require multiple steps or diverse knowledge. They are often employed in scenarios that demand role-playing [92] or coordinated actions [148], showcasing their versatility in handling intricate, multi-faceted challenges [149].
While these prompting techniques share some commonalities, they have distinct characteristics and applications. The Zero-shot Prompt method guides the model using a sentence or paragraph as a prompt text. The other three methods build upon this with various modifications and expansions. The Few-shot Prompt method adds task-solving examples to the prompt texts to assist in task completion. The CoT Prompt method employs multi-step chain thinking, using different prompt texts to complete the task at each step. CoT can also be combined with the few-shot method, using CoT’s the multi-step thinking process of CoT as examples to guide the generation of the thought chain. The agent method differs from the previous three methods in adopting a collaborative approach with multiple LLMs. Each LLM can use any prompt form, such as zero-shot, few-shot, or CoT, and the information output by each LLM is combined to complete the task. Figure 3 illustrates the similarities and differences among the four methods.
@@ 5.1.1 Zero-shot for Affective Understanding
In the field of affective understanding, zero-shot prompting is initially applied. Researchers use descriptions of sentiment analysis tasks as prompts, placing them before the target text to guide LLMs in generating sentiment for the target text. Through this method, researchers can observe the resulting sentiment analysis outcomes, thereby gaining a more comprehensive understanding of the performance and potential of various LLMs in AU [45, 150, 151, 42].
To enhance the performance of LLMs in AU, researchers develop several advanced prompting techniques that go beyond simple task descriptions. One such approach is a role-playing zero-shot prompting method [40], where LLMs are assigned a specific role, such as an emotion expert, to perform sentiment analysis tasks. By assigning roles, the emotional capabilities of the model can be better utilized. Additionally, a straightforward yet effective approach is called EmotionPrompt, a series of emotion stimulus questions based on psychology are incorporated into the original task prompts [63] to explore the sentiment analysis capabilities of LLMs by embedding specific information in the prompts. In Speech Emotion Recognition (SER), researchers incorporates information from acoustic feature descriptors into the prompts [152], combining numerical acoustic features into the prompts to identify the underlying emotions conveyed in the speech.
Although these zero-shot prompting methods using fixed prompt content achieve notable performance, their flexibility and diversity are limited due to the constraints of the prompt content. Therefore, in the field of multimodal sentiment analysis, researchers explore the use of LLMs to generate prompts based on specific needs [39], and then use these generated prompts as inputs to guide LLMs in analyzing the sentiments involved in image-text pairs.
Additionally, more attention must be paid to interpretation or emotional response in a range of subjective tasks such as Metaphor Recognition [153], and Dark Humor Detection [52]. RiC (Reasoning in Conversation) [154] is a Zero-shot approach that addresses these subjective tasks through dialogue simulation. RiC extracts useful contextual information by simulating conversations, providing the underlying knowledge behind the dialogue for the final answer.
@@ 5.1.2 Few-shot for Affective Understanding
Few-shot prompting is introduced to handle more challenging tasks, such as complex Aspect-Based Sentiment Analysis (ABSA) and Multimodal Sentiment Analysis (MSA).
For the challenging Aspect-Sentiment Quad Prediction (ASQP) [155] task, the Self-Consistent Reasoning-based Aspect-Sentiment Quad Prediction model, named SCRAP, is proposed [44]. This model uses prompt examples to generate reasonable reasoning-related content, known as few-shot CoT prompting, for subsequent fine-tuning. By leveraging a few reasoning strategies as prompts to enhance LLM reasoning and integrating the generated reasoning into the subtasks of ABSA, the ability of the model to handle complex reasoning tasks is improved, along with its interpretability and accuracy.
Few-shot Prompt method is also applied to MSA in AU. WisdoM is proposed to address four Multimodal Entity-Based Sentiment Analysis (MEBSA) subtasks [156]. WisdoM converts visual inputs into textual counterparts regarded as auxiliary sentences and selects positive and negative instances based on text similarity to construct examples. In the context generation stage, LVLMs [157, 158] use the prompt templates generated in the prompt templates generation stage to generate context that explicitly contains world knowledge based on a given image and sentence demonstration. Although WisdoM is a straightforward and effective plugin framework that significantly enhances the performance of multimodal sentiment analysis, the rich information in visual content still need to be fully utilized. Therefore, compared to the easy-to-understand semantic information from the text modality, fully extracting and utilizing the content from the image modality with multimodal LLMs remains a significant challenge.
@@
Additionally, in the Depression Detection task, Few-shot Prompt method guides data generation to enhance the Structural Element Graph (SEGA) [159] further. Researchers use explicit principles to guide the synthesis process and provided LLM with data generation examples to obtain high-quality synthetic data.
5.1.3 CoT for Affective Understanding
The information provided by zero-shot or few-shot prompting through a single input-output process must be more sufficient to support the completion of complex AU tasks. Therefore, researchers explore the CoT method, which incorporates a series of intermediate reasoning steps during the prompting process to enhance the model’s understanding of inputs, thereby better grasping the emotions and nuanced meanings.
Due to the complexity of affective understanding, introducing information that aids in affective understanding can significantly enhance the model’s ability to perceive and process emotions. Therefore, researchers propose a three-hop reasoning chain framework called THOR [46] for Implicit Sentiment Analysis (ISA). Based on the importance of hidden opinions in conversations, the THOR CoT framework sequentially extracts fine-grained aspects, analyzes the latent opinions on these aspects, and derives accurate implicit sentiments based on this information.
Unlike conventional methods that directly extract regular information from known texts, the Chain of Empathy (CoE) [56] adopts a psychological perspective, incorporating various psychotherapy methods such as Cognitive Behavioral Therapy (CBT) [160], Dialectical Behavior Therapy (DBT) [161], Person-Centered Therapy (PCT) [162], and Reality Therapy (RT) [163]. It designs four thinking chain methods to interpret user’s psychological states of users from a professional standpoint. Ultimately, CoE uses the information from these psychotherapy methods to infer human emotional states,achieving a more accurate understanding and processing of emotion. Drawing inspiration from the Stimulus-Appraisal-Emotion process within the cognitive appraisal theory of emotion generation, the ECR-Chain framework [51] undertakes a multi-step approach. First, the ECR-Chain summarizes the dialogue topics. Next, it generates speaker responses to these specific topics using LLM. Subsequently, the model infers the user’s target appraisal and extracts stimulus information from the dialogue. Finally, based on this information, the ECR-Chain infers the underlying cause of the emotion.
Methods such as COE and ECR-Chain generate beneficial inference information based on task requirements and achieve good results. However, they face significant challenges in complex extraction tasks, such as Joint Aspect Sentiment Analysis (JASA) and Emotion-Cause Pair Extraction (ECPE). A multi-step reasoning approach is proposed to decompose the tasks. A multi-step reasoning method is proposed to decompose the task. Specifically, to improve the performance of sentiment analysis (SA), RP-CoT [40] first identifies the sentiment polarity information corresponding to various aspects of the text instance and then classifies the overall sentiment polarity of the entire text based on the obtained fine-grained sentiment information. Similarly, when performing the task of ECPE, DECC [50] first identifies the emotions present in the given context and then locates and reserves the specific utterances (emotional clauses) associated with each emotion. Finally, for each emotion, the most probable clause is selected as its cause, and the pairs of emotional and cause clauses are outputted.
In addition to using CoT to generate information that aids in understanding emotions and decomposing complex tasks, the CoT method is also applied to prompt construction. The Evolutionary Multi-Objective (EMO) process [164] inputs manually crafted original prompts into LLMs and evaluates their fitness through generated stories. EMO then randomly selects two prompts from the original set and recombines them using crossover operators. Applying a new set of mutation operators to the recombined prompts, it ultimately produces new prompts that are better suited to the task.
The CoT Prompt method can also be combined with Few-shot Prompt method. In the task of Market Sentiment Analysis [165], it is necessary to master financial market and social media terminology, making manual evaluation extremely challenging. The lack of high-quality labeled data limits the application of traditional supervised learning methods. Researchers have drawn on the CoT concept by constructing contextual examples and using Few-shot Prompt method to prompt LLMs to generate CoT summaries [166]. This guides LLMs to recall relevant financial domain knowledge beforeconclusions, thereby injecting relevant knowledge into Market Sentiment Analysis.
@@ 5.1.4 Agent for Affective Understanding
In previous studies, the standard paradigm for affective understanding relies on a single LLM and makes decisions [46, 156]. However, this method has a critical drawback: the output generated by a single LLM may not provide the optimal decision. Therefore, the Agent technology is further proposed to enable the collaboration of multiple LLMs.
The Generator-Discriminator Role-switching Decision-Making framework is proposed to achieve multi-LLM negotiation for sentiment analysis [41]. This framework consists of a reasoning-infused generator, responsible for providing decisions and rationale, and an explanation-deriving discriminator, which evaluates the credibility of the generator. The generator and discriminator iteratively communicate until a consensus is reached. By leveraging the complementary capabilities of the two LLMs, sentiment interpretation and correction can be better achieved. The PANAS framework [167] is designed with multiple agents to perceive different information. New natural language experiences are input into Agent 1, which summarizes and extracts insights from memories to create norms. Agent 2 conducts context understanding, highlighting the differences between the current situation and the norms. Finally, Agent 3 evaluates emotional responses based on the provided context understanding.
In the field of Financial Sentiment Analysis (FSA), HAD [168] summarizes five common types of errors in FSA tasks and instantiates five specialized agents based on this prior knowledge: the mood agent, the rhetoric agent, the dependency agent, the aspect agent, and the reference agent. Each agent focuses on a specific type of error that LLMs are prone to in given FSA tasks. Ultimately, the FSA results are obtained through a shared discussion that considers the outputs of all agents.
@@ 5.2 Affective Generation
Prompt engineering has also been introduced in the field of Affective Generation to improve performance across various tasks and scenarios. The following introduces four applications of prompt engineering techniques in Affective Generation.
@@ 5.2.1 Zero-shot for Affective Generation
To adapt prompts to the task characteristics of Affective Generation, researchers propose the Perspective-Taking method [169], which enhances the capability of LLMs to generate empathetic dialogues by enabling perspective shifts. Within the HEF framework [54], small empathy models are first trained to generate the emotion cause and the main emotional categories in the dialogue, which are then provided as input to the LLM. This approach gives the model more accurate information, resulting in more precise empathetic responses generated under zero-shot prompts. Researchers design the TCG pipeline [62] based on GPT-3.5-turbo for the task of opinion summarization. Through sentence opinion extraction, grouping, and clustering, the pipeline ultimately produces a summary of the long-form’s opinions.
In the emotion-conditioned text generation task, researchers optimizes prompts by iteratively adding, replacing, and deleting the original prompts [170]. The optimized prompts better achieve the expected emotions, resulting in improved emotional text generation. Since the output of LLMs largely depends on prompts, users needing more background knowledge might experience a diminished interaction with chatbots due to the absence of contextual understanding. To address this issue, PromptMind [171] generates multiple contextually relevant prompts during the conversation for users to choose from, thereby prompting LLMs to produce answers that better meet user expectations and enhance human-chatbot interaction.
@@ 5.2.2 Few-shot for Affective Generation
In Sibyl framework [172], researchers randomly select a sample as a demonstration to prompt the LLM to generate four categories of visionary commonsense related to the dialogue as data for subsequent training of the dialogue generation model. Sibyl improves the empathy generation ability of the LLM [53]. In mixed-initiative dialogue generation tasks like PersuasionForGood and Emotional Support Conversations, generating responses by setting different strategies as examples is also a way to use few-shot prompting in generation tasks [173].
Researchers indicate that generated responses utilizing prompt engineering demonstrate high quality and adherence to semantic controls [173]. However, in emotional support tasks, the inherent preference of LLMs for specific strategies can hinder effective support despite improving the robustness of predicting appropriate strategies. To address this, researchers incorporate strategy descriptions and randomly select 2-Shot examples into prompts to enhance the LLM’s understanding of each strategy [174]. LLaMA2-7B [14] is leveraged for planning the subsequent action, followed by LLM generation of responses aligned with this strategy. Finally, the number of examples within the prompt is expanded. Notably, results suggest that while providing an appropriate number of examples can enhance preference bias, simply increasing the number of examples may only sometimes yield better results.

### v2

@@ 4.4 Challenge
  • Despite their known universality, LLMs often struggle to simultaneously excel in both AU and AG within the domain of AC. It remains a significant challenge to develop a unified model capable of effectively addressing AU and AG.
@@ 5 Prompt Engineering
Prompt engineering Brown et al. (2020) guides LLMs to produce desired outputs by designing appropriate prompts, effectively improving the accuracy and reliability of AC tasks. This section reviews recent advances in prompt engineering for AC, focusing on four techniques: zero-shot prompting, few-shot prompting, Chain-of-Thought (CoT) prompting, and agent.
Zero-shot Prompting Brown et al. (2020) generates text for a specific topic or domain without any demonstrations. It is simple to use, highly adaptable, and especially suitable for scenarios where data are scarce or hard to obtain.
Few-shot Prompting Brown et al. (2020) supplies the model with a small set of task-relevant examples or demonstrations. These examples, together with the task description, are used as prompts to help the model better understand the task, reason more effectively, and produce high-quality outputs.
Figure 4: The overview of prompt engineering.
Chain-of-Thought Wei et al. (2022b) adopts a multi-step reasoning paradigm that guides the LLMs to reason logically and explicitly reveal the intermediate steps of its reasoning chain. Chain-of-Thought enables LLMs to better leverage context, produce more accurate and interpretable answers, and perform better on complex reasoning tasks.
Agent Park et al. (2023) assigns clear roles, objectives, and collaboration protocols to multiple LLMs, driving them to collaborate via message passing and feedback loops to complete tasks. Agent effectively handles task decomposition and planning, tool invocation, and memory management, and excels in open-domain, multi-stage, highly interactive scenarios.
While these prompting paradigms share commonalities, they differ in emphasis and application contexts. Zero-shot prompting directs the model with task instructions alone, without examples. Few-shot prompting appends a small number of exemplars to the instruction to align format and decision boundaries. Chain-of-thought prompting explicitly exposes intermediate reasoning steps or uses trigger phrases to elicit step-by-step thinking. Agent methods orchestrate a workflow of one or more agents and tools to support task decomposition, planning, memory, and self-verification, and to aggregate multiple outputs. Figure 4 summarizes the distinctions and composability of the four approaches.
@@ 5.1.1 Zero-shot for Affective Understanding
In affective understanding (AU), zero-shot prompting has been applied. Researchers preset a task description as the prompt to the target text to guide LLMs to produce sentiment labels Amin et al. (2023b); Ouyang et al. (2024); Al Asad et al. (2023); Stigall et al. (2024). Building on this, incorporating dialogue context Hong et al. (2025), multi-label prompts Zhou et al. (2023a), constraint-based prompts Bai et al. (2024), and task-specific prefixes Cabello & Akujuobi (2024) further improves LLM performance on AU tasks. Researchers explore a variety of zero-shot prompting strategies to improve the performance of LLM-based models on the AU tasks. LLMs are cast as emotion experts Wang & Luo (2023), assigned specific personas Niu et al. (2024), or given gendered roles Plaza-del Arco et al. (2024) to perform sentiment analysis. Psychologically inspired prompts are also introduced, such as emotional stimulation questions Li et al. (2023a) and integrating narrative style elements extracted by HEART Shen et al. (2024). In Speech Emotion Recognition (SER), acoustic features are verbalized into textual descriptors Santoso et al. (2024), prompts are constructed in natural language Wu et al. (2024d), and emotional cues in speech are distilled into prompt prefixes Liang et al. (2024), thereby further improving recognition accuracy. Although fixed-template zero-shot prompting can be effective, its fixed content limits flexibility and diversity. Accordingly, multimodal sentiment analysis explores having LLMs generate task-specific prompts on demand and then using these prompts to guide the analysis of emotions in image–text pairs Wang et al. (2024f).
Additionally, subjective tasks such as Metaphor Recognition Mohler et al. (2016) and Dark Humor Detection Hoffmann et al. (2022) require greater attention to emotional responses. Reasoning in Conversation (RiC) Wang et al. (2024g) is a zero-shot paradigm that addresses tasks via dialogue simulation, extracting key information to support the final answer.
@@ 5.1.2 Few-shot for Affective Understanding
The pivotal role of demonstrations in prompt drisves exploration of the few-shot prompting. Early methods required manually curating and inserting demonstrations Zhou et al. (2024). With a range of selection strategies emerging Zhang et al. (2024d), a central challenge is how to more effectively elicit and unlock LLMs’ few-shot capabilities for sentiment analysis. For the challenging Aspect-Sentiment Quad Prediction (ASQP) task Zhang et al. (2021), The Self-Consistent Reasoning-based Aspect-Sentiment Quad Prediction model (SCRAP) Kim et al. (2024) integrates pre-generated reasoning demonstrations into the task’s reasoning process to strengthen LLM reasoning. Few-shot prompting has also been introduced into multimodal sentiment analysis (MSA) for AU. WisdoM Yang et al. (2024a) targets four multimodal entity-based sentiment analysis (MEBSA) subtasks by converting visual inputs into textual surrogates treated as auxiliary statements, and constructing demonstrations by selecting positive and negative instances based on textual similarity. During context generation, LVLMs Liu et al. (2024a); Ye et al. (2024) use prompt templates, together with the given image–sentence demonstrations, to produce contexts that explicitly incorporate world knowledge.
@@
Additionally, in Depression Detection task, few-shot prompting guides data generation to enhancing the Structural Element Graph (SEGA) Chen et al. (2024c). Researchers follow principles to steer the synthesis process and provide the LLM with generation exemplars, yielding high-quality synthetic data.
5.1.3 Chain-of-Thought for Affective Understanding
Zero-shot or few-shot prompting carries limited information in a single-turn interaction, making it insufficient for complex AU tasks. To address this, researchers introduce CoT Lee et al. (2024), which explicitly adds intermediate reasoning steps to better handle complex tasks. Given the complexity of AU, THOR Fei et al. (2023) employs a three-hop CoT framework for Implicit Sentiment Analysis (ISA). It sequentially extracts fine-grained aspects, analyzes latent opinions, and infers implicit sentiment. In addition, injecting causal relations as causal cues into the model Lyu et al. (2024) can enhance LLMs’ sentiment analysis performance. Unlike traditional methods that extract regularities directly from text, Chain of Empathy (CoE) Lee et al. (2023b) combines multiple psychotherapeutic approaches Beck (1979); Linehan (1987); Cooper & McLeod (2011); Wubbolding et al. (2017), designs four types of chains of thought, and interprets users’ psychological states within a professional framework, thereby enabling a more precise understanding and handling of emotions. Meanwhile, other work Manzoor et al. (2024) adopts a psychological information–mining perspective and explores LLMs’ potential for empathetic understanding. ECR-Chain Huang et al. (2024) draws on appraisal theory’s “stimulus–appraisal–emotion” process and employs multi-step reasoning to infer the deeper causes of emotions.
Methods such as CoE and ECR-Chain generate task-specific reasoning and achieve good results, but they still face challenges on complex extraction tasks such as Joint Aspect Sentiment Analysis (JASA) and Emotion-Cause Pair Extraction (ECPE). Accordingly, RP-CoT Wang & Luo (2023) adopts multi-step reasoning to decompose the task: it first identifies aspect-level sentiment polarities and then determines the overall polarity. Likewise, DECC Wu et al. (2024b) first detects emotions and locates the corresponding emotion clauses, and then, for each emotion, selects the most probable cause clause and outputs emotion–cause pairs. Beyond generating intermediate reasoning for emotion understanding and task decomposition, CoT also supports prompt construction. The Evolutionary Multi-Objective (EMO) process Baumann & Kramer (2024) starts from human-written seed prompts, feeds them to an LLM and evaluates fitness via the generated stories. It then performs crossover recombination and applies mutation operators, iteratively producing new prompts.
The CoT can be combined with few-shot prompting. For Market Sentiment Analysis Chen et al. (2018), the complexity of financial and social-media terminology and data scarcity constrain traditional supervised methods. Researchers construct contextual exemplars and use few-shot prompts to elicit CoT summaries Deng et al. (2023a), prompting the model to retrieve and integrate domain knowledge before concluding and thereby injecting relevant financial knowledge into the Market Sentiment Analysis.
@@ 5.1.4 Agent for Affective Understanding
Prior studies mostly adopt a single LLM decision paradigm Fei et al. (2023); Yang et al. (2024a). However, a single model’s output does not reliably approach the optimum. To address this, Agent techniques orchestrate collaboration and evaluation among multiple LLMs, enabling more robust ensemble decisions.
To achieve multi-LLM negotiation in sentiment analysis, the Generator-Discriminator Role-switching Decision-Making framework Sun et al. (2023) is proposed. The generator injects reasoning and provides decisions with rationales, the discriminator explains and assesses credibility, and the two iterate until they reach consensus, leveraging model complementarity for more accurate emotion interpretation and correction. Building on this idea, agent design expands with an automated 5W1H question-generation module to predict emotions from social information Wei et al. (2024). Multi-agent interaction also builds an emotion contagion model that characterizes how emotions propagate within groups and their effects Van Haeringen et al. (2023). The PANAS framework Regan et al. (2024) employs three specialized agents for memory, context understanding, and evaluation. Meanwhile, AnnaAgent Wang et al. (2025a) uses multi-session memory to capture users’ long-term emotional dynamics, offering a practical route to more realistic and deeper digital companions.
In Financial Sentiment Analysis (FSA), HAD Xing (2024) summarizes five common error types and, accordingly, instantiates five specialized agents—sentiment, rhetoric, dependency, aspect, and reference. Each agent targets the error-prone dimension for LLMs, and their outputs are aggregated and deliberated to produce the final FSA decision.
@@ 5.2 Affective Generation
Prompt engineering is also used to improve the performance of AG. The following introduces the application of four prompt engineering technologies in AG.
@@ 5.2.1 Zero-shot for Affective Generation
To align prompts with the characteristics of AG, researchers propose Perspective-Taking Lee et al. (2023a), which enhances LLMs’ empathetic dialogue generation by shifting viewpoints. In HEF Yang et al. (2024c), a small empathy model first extracts emotion causes and primary emotion categories. These are then provided to the LLM under zero-shot prompting to produce more precise empathetic responses. The TCG pipeline Bhaskar et al. (2022), built on GPT-3.5-turbo, performs sentence-level opinion extraction, grouping, and clustering to generate long-form summaries of affective opinions.
In the emotion-conditioned text generation task, researchers optimizes prompts by iteratively adding, replacing, and deleting the original prompts Resendiz & Klinger (2023). The optimized prompts better achieve the expected emotions, resulting in improved emotional text generation. Since the output of LLMs largely depends on prompts, users needing more background knowledge might experience a diminished interaction with chatbots due to the absence of contextual understanding. To address this issue, PromptMind Su et al. (2023) generates multiple contextually relevant prompts during the conversation for users to choose from, thereby prompting LLMs to produce answers that better meet user expectations and enhance human-chatbot interaction.
@@ 5.2.2 Few-shot for Affective Generation
In the Sibyl framework Wang et al. (2023c), researchers randomly select a single example as a demonstration to prompt the LLM to generate four types of prospective commonsense related to the dialogue, which are then used as training data for subsequent dialogue-generation models. For mixed-initiative dialogue generation tasks such as PersuasionForGood and Emotional Support Conversations, responses are generated conditioned on different exemplar strategies Chen et al. (2023a).
Researchers indicate that generated responses utilizing prompt engineering demonstrate high quality and adherence to semantic controls Chen et al. (2023a). However, in emotional support tasks, the inherent preference of LLMs for specific strategies can hinder effective support despite improving the robustness of predicting appropriate strategies Kang et al. (2024). To address this issue, researchers conduct extensive experiments involving Self-contact approaches, external strategy planning, and example expansion, demonstrating that External-contact approaches help reduce LLM preference bias and enhance the performance of emotional support.